[Hallc_running] [New Logentry] RC Update

latif at jlab.org latif at jlab.org
Mon Nov 12 11:35:01 EST 2018

Logentry Text:
- CSV kinematic group-5 was completed during the last owl shift.

- SHMS polarity was changed to switch to kinematic group-6 (relevant entry: https://logbooks.jlab.org/entry/3624805). Setting 6-3 was completed before beam was taken away at 7:30 AM.

- DAQ start-of-run auto-log was found missing in some cases. Otherwise, smooth running.

- The accelerator was able to provide 50 uA in some cases. In some cases it limited us to 20-30 uA. Hall-C higher current is part of their beam studies tomorrow.

- Currently, the accelerator is down for system recovery which is scheduled to last ~8 hours. Nobody requested any access to the hall.

- The accelerator will perform extended (~8 hours) beam studies with high current starting at 8:30 AM tomorrow. Hall-C will remain locked up.

We will meet today at 4:30 PM in the counting house (second floor) to discuss the update and run plan.
Bluejeans link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3624981
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181112/8f00ff51/attachment.html>

More information about the Hallc_running mailing list